# The EU AI Act, Stakeholder Needs, and Explainable AI: Aligning Regulatory Compliance in a Clinical Decision Support System

### v1

@@ Abstract
Explainable AI (XAI) is a promising solution to ensure compliance with the EU AI Act, the first multi-national regulation for AI. XAI aims to enhance transparency and human oversight of AI systems, particularly “black-box models”, which are criticized as incomprehensible. However, the discourse around the main stakeholders in the AI Act and XAI appears disconnected. While XAI prioritizes the end user’s needs as the primary goal, the AI Act focuses on the obligations of the provider and deployer of the AI system. We aim to bridge this divide and provide guidance on how these two worlds are related. By fostering an interdisciplinary discussion in a cross-functional team with XAI, AI Act, legal, and requirements engineering experts, we walk through the steps necessary to analyze an AI-based clinical decision support system to clarify the end-user needs and assess AI Act applicability. By analyzing our justified understanding using an AI system under development as a case, we show that XAI techniques can fill a gap between stakeholder needs and the requirements of the AI Act. We look at the similarities and contrasts between the legal requirements and the needs of stakeholders. In doing so, we encourage researchers and practitioners from the XAI community to reflect on their role towards the AI Act by achieving a mutual understanding of the implications of XAI and the AI Act within different disciplines.
@@ 1 Introduction
After long and complicated negotiations between the European Parliament, the EU Commission and the EU Member States, the AI Act entered into force in August 2024 (European Parliament and Council, 2024). AI-based systems are now regulated under a comprehensive set of product safety rules for the first time. In the coming years, different parts of the EU AI Act will become applicable. Providers and deployers of AI systems need to learn how to apply the different articles of the act, while notified bodies and certification authorities will need to establish a process to assess these systems according to the regulations. Many standardization efforts are already underway, but there will be significant uncertainty about how the AI Act has to be applied and how it relates to different aspects of AI currently under research.
Regulations such as the AI Act are especially important in high-risk artificial intelligence (AI) systems to ensure that they are safe and respect fundamental rights and values (Van Kolfschooten and Van Oirschot, 2024). The AI Act follows a risk-based approach: the higher the risk, the stricter the rule. Especially in high-risk systems, it describes many obligations for the provider and the deployer of AI systems.
@@
cdss, which may fall under the definition of a high-risk AI system in the EU AI Act, place the responsibility for the high-risk decision on the end-user (typically, clinicians). However, the AI Act does not directly address end-user needs and instead primarily focuses on regulating the providers and deployers of AI systems. One of the few mentions of end-users is the statement that high-risk AI systems must be designed so that humans can effectively oversee these systems. But there are no obligations in terms of technical requirements that put the user’s needs at the center of its goals (Van Kolfschooten and Van Oirschot, 2024).
In contrast, nowadays more explainable artificial intelligence (XAI) researchers suggest putting end-users and their needs at the center of attention (Langer et al., 2021). XAI methods provide descriptions of the outputs of opaque AI models, thereby satisfying the needs of the end-users. This is very helpful in high-risk applications such as healthcare. Previous work has started exploring the role that XAI plays w.r.t. ensuring the trustworthiness of high-risk applications under the EU AI Act (Panigutti et al., 2023; Kiseleva et al., 2022; Longo et al., 2024; Esmaeilzadeh, 2024). XAI can contribute to implementing transparency and human oversight, which are key principles of the AI Act (Panigutti et al., 2023). Taking this line of thinking further, we believe that XAI—if desigend correctly—could play a crucial role in fulfilling the obligations of deployers and providers of AI systems. Therefore, we ask:
@@
Research on the AI Act and research on XAI are often conducted independently, lacking interdisciplinary collaboration (Gyevnar et al., 2023). By providing a joint understanding of the role of XAI in the AI Act from an interdisciplinary perspective representing various stakeholder needs, we…
@@
    provide an analysis of the implications of the AI Act in relation to adopting AI in healthcare (see Section 4)
@@
    describe how the needs of the CDSS stakeholders can be captured through XAI (see Section 5)
@@
    assess the similarities and contrasts between the legal requirements of the AI Act and the stakeholder desiderata of XAI (see Section 6).
With this, our work is the first that analyzes the role of XAI in the AI Act in a cross-functional team, on a real-world CDSS that takes user-centered needs into account and contributes to the state-of-art by mapping obligations of the EU AI Act with the obligations of stakeholders and how to fulfill them using XAI. With this, we encourage researchers and practitioners from the XAI community to reflect on their role towards the AI Act by bridging the gap towards end-users.
@@ 2 Related Work and Background
In this section, we give an overview of the AI Act and XAI. We also discuss relevant related work. A number of studies already addressed the question of how AI can be applied in healthcare, using ethical and legal considerations (Aboy et al., 2024; Porsdam Mann et al., 2024; Karimian et al., 2022; Kiseleva et al., 2022). The function of XAI in fostering transparency and human oversight has also been the subject of considerable debate (Panigutti et al., 2023). Other works have dealt with the role of XAI in the AI Act (e.g., (Gyevnar et al., 2023)) and have given an overview of the AI Act and XAI (e.g., (Panigutti et al., 2023)). Nevertheless, it remains unexplored how XAI can be effectively utilized to fulfill the obligations of stakeholders as defined in the AI Act.
@@ 2.1 Background: The EU AI Act
In 2024 the European Union ratified the regulation of AI as technology (European Parliament and Council, 2024). The new regulation was motivated by a need to create trust for AI with respect to product safety, health and fundamental rights. The AI Act thus incorporates AI systems into the EU’s existing framework for product safety. Certain AI systems will need to undergo conformity assessment and subsequently carry the CE-mark as a sign of being a certified product, just like elevators, machinery and medical devices (see Section 2.1).
The regulation also gives authorities an important role in guaranteeing the responsible development and usage of AI, as public sector bodies are to conduct market surveillance but also foster AI innovation (see e.g., (Burden and Stenberg, 2022)). There are two central actors from the perspective of the regulation: the provider (see Section 2.1) and the deployer (see Section 2.1) of an AI system. These actors are obliged to take on responsibility for product safety.
@@
The conformity assessment for CE-marking AI systems can then be done in two ways. The first is through self-assessment where the provider of the system certifies that the requirements are met. This is also referred to as internal conformity assessment. The second conformity assessment procedure involves a notified body, an organization that has been accredited to be competent in conformity assessment of certain products. The role of the notified body is to give a neutral assessment to what extent the requirements are met. The nature of the AI system will determine which process is to be applied (Burden et al., 2024).
@@
There will be harmonized standards issued by the European standardization organizations CEN and CENELEC that further details the requirements in Art. 9 to 15 and 17. An overview of the standardization effort can be found on the two organizations’ joint technical committee’s website (European Committee for Standardization, 2025).
@@ 2.2 Background: XAI Terminology
As indicated, current research on explainable artificial intelligence (XAI) provides a new perspective on the relationship between an AI system and their affected parties. Therefore, we provide a brief nomenclature of relevant key concepts and terms. We define the key terms related to XAI, examine the various stakeholder terminologies, and clarify a common terminology of XAI methods.
@@
The terms explainability, interpretability, and transparency are often used interchangeably (Al-Ansari et al., 2024; Renftle et al., 2024; Hoffman et al., 2023; Barredo Arrieta et al., 2020; Mohseni et al., 2021). We distinguish these terms by categorizing them into approaches like Explainable AI and Interpretable AI, and desired properties such as Interpretability, Transparency, and Explainability (Mohseni et al., 2021). Explainable AI provides human-understandable explanations for predictions made by complex AI systems that are not inherently interpretable (Wang et al., 2019). Interpretable AI, on the other hand, requires no additional methods for human understanding, as these systems are inherently interpretable (Mohseni et al., 2021). Interpretability refers to the ability to explain or provide meaning in understandable terms, while Transparency indicates a model is interpretable by design (Barredo Arrieta et al., 2020).
@@
XAI goals such as trustworthiness, informativeness, and fairness have been widely studied (Al-Ansari et al., 2024; Renftle et al., 2024; Hoffman et al., 2023; Barredo Arrieta et al., 2020; Mohseni et al., 2021). Recently, these goals have shifted to focus more on specific stakeholders (Samek and Müller, 2019; Biecek and Samek, 2024), their desiderata (Langer et al., 2021), and expertise levels (Mohseni et al., 2021). Stakeholders’ interests, goals, expectations, needs, and demands regarding AI systems are referred to as their desiderata (Langer et al., 2021).
Understanding different stakeholders in XAI is essential for creating systems that meet various desiderata (Langer et al., 2021). Different groups have distinct interests, and recognizing these differences is key to developing methods that align with their specific desiderata. The success of an XAI method relies on satisfying stakeholders’ desiderata, which drives explainability approaches that enhance understanding. Langer et al. (2021), following Barredo Arrieta et al. (2020), categorize stakeholders into five classes: users, (system) developers, affected parties, deployers, and regulators. This classification ensures all relevant needs are considered, with the challenge being to provide suitable explanations for each group (Mohseni et al., 2021).
Following Mohseni et al. (2021), end-users are divided into three groups based on their expertise level: Novice Users, who use AI systems daily with little machine learning (ML) knowledge; Data Experts, including data scientists and domain experts using ML for analysis; and AI Experts, comprising ML scientists and engineers designing and implementing ML algorithms and XAI methods.
@@
We utilize common XAI taxonomies relevant to our CDSS system, following Speith (2022).
@@
Scope refers to local explanations that detail specific input-output relations, while global explanations describe the overall ML model. Examples include model visualizations (Liu et al., 2017) and decision rules (Lakkaraju et al., 2016).
Other taxonomies include input data (e.g., image, time-series), output format (e.g., numerical, visual), and problem type (e.g., classification, regression).
@@ 2.3 Related Work
As previously mentioned, there are several studies that address the question of how AI can be applied in healthcare in general. Table 1 provides an overview of the related work papers and indicates how we position our own contribution.
@@
Paper AI Act XAI Healthcare Real World User-Centric
Gyevnar et al. (2023)
Pavlidis (2024)
Fresz et al. (2024)
Onitiu (2023)
Solaiman and Malik (2024)
Van Kolfschooten and Van Oirschot (2024)
Panigutti et al. (2023)
@@
Gyevnar et al. (2023) explore the transparency gap between the AI Act and XAI understanding, emphasizing the differing conceptualizations. They focus on stakeholder engagement and how XAI can address these concerns, aligning with our approach. Identified solutions include concentrating on legal definitions, AI system output, and data-related requirements to meet transparency standards. It’s important to note that transparency requirements are broad and should not be seen as concrete engineering instructions. XAI contributes to transparency but should not be viewed as a standalone solution, instead facilitating transparency while adhering to its foundational principles.
@@
Pavlidis (2024) explores the role of XAI as a tool to bridge the gap between the opaque nature of ML decisions and the human demand for comprehensibility. In essence, XAI serves as a mechanism to align the objectives of the AI Act, such as enhancing transparency, with the preferences of end-users. The author examines the integration of XAI into the AI Act, with a focus on critical issues like standard setting, oversight, and enforcement. Fresz et al. (2024) review legal requirements for explainability of AI systems from the AI Act. Furthermore, they describe legal bases for the use of XAI and derive requirements the XAI methods need to fulfill. All three articles deal with the role of XAI in the AI Act. Our work extends this view by elaborating the role of XAI in a real-world healthcare scenario and by having a incorporating the the end-user needs.
@@
A more domain-specific approach is presented by Onitiu (2023), who aims to translate technical safeguards in Art. 13-14 of the AI Act proposal for the use of medical diagnostic systems. The author proposes a broader perspective on the socio-legal implications of technical safeguards, such as post-hoc explainability for decision-making. The author emphasizes that post-hoc XAI methods must focus on harmonizing human expertise with patient-oriented values.
@@
Similarly, Solaiman and Malik (2024) explore the AI Act in the healthcare domain with a user-centric focus on AI. Hereby, XAI plays a subordinate role. Instead, the authors focus on the doctor-patient relationship that evolved from a relation where the doctors held the dominant role and the patient follows their guidance without explanation, to a relationship where the patient is seen as active participants in their healthcare decisions. AI and especially XAI can have an important impact providing more information to both clinicians and patients, thus enhancing patient autonomy.
@@
Van Kolfschooten and Van Oirschot (2024) analyze the implications of the AI Act for the healthcare domain. Here, they give an overview of the obligations for various healthcare stakeholders. Then they use their insights to provide a guideline to address the unique needs of the stakeholders and make recommendations for the implementation and standardization of the act. They also highlight the importance of collaborating with different stakeholders like AI developers, healthcare professionals, patient communities, and regulators to develop those guidelines.
@@
In a more application-oriented study, Panigutti et al. (2023) conduct an interdisciplinary analysis of the AI Act’s approach to opaque AI systems, aiming to clarify XAI’s role. They focus on Art. 13, concerning transparency, and Art. 14, regarding human oversight, providing an overview of XAI terminology and key concepts. The authors also review the XAI field, noting challenges and limitations consistent with previous findings. They analyze how the AI Act addresses AI system opacity through transparency and human oversight, presenting a use case involving an AI-based proctoring system without XAI tools. In contrast, our study extends beyond Art. 13 and 14 and examines other articles that establish obligations for healthcare stakeholders.

### v2

@@ Abstract
Explainable AI (XAI) is a promising route to comply with the EU AI Act, the first multinational AI regulation. XAI enhances transparency and human oversight of AI systems, especially ”black-box“ models criticized as incomprehensible. Yet discourse about the AI Act’s stakeholders and XAI remains disconnected: XAI increasingly prioritizes end users’ needs, while the AI Act focuses on providers’ and deployers’ obligations. We aim to bridge this divide and offer practical guidance on their relationship. Through interdisciplinary discussion in a cross functional team of XAI, AI Act, legal, and requirements-engineering experts, we outline steps to analyze an AI-based clinical decision support system, clarify end-user needs, and assess AI Act applicability. Using an AI system under development as a case study, we show how XAI techniques can help reconcile stakeholder needs with AI Act requirements and fill gaps between usability and regulatory demands. We compare similarities and differences between legal obligations and end-user needs, identify tensions, and point to concrete design choices and trade-offs. We invite researchers and practitioners in XAI to reflect on their role relative to the AI Act and to develop mutual understanding across disciplines. While XAI can help implement core AI Act principles such as transparency and human oversight, it should be considered one element of a broader compliance strategy that also requires standardization, legal interpretation, documentation, organizational processes, governance, testing, and ongoing monitoring and auditing practices. Our findings yield actionable recommendations for integrating XAI into product development, compliance workflows, and stakeholder communication, informing policy-making and standards development.
Keywords European Artificial Intelligence Act, Explainable Artificial Intelligence, Clinical Decision Support, Healthcare, Interdisciplinary Research
@@ 1 Introduction
After long and complicated negotiations between the European Parliament, the EU Commission, and the EU Member States, the AI Act entered into force in August 2024 (aiact). AI-based systems are now regulated under a comprehensive set of product safety rules for the first time. In the coming years, different parts of the EU AI Act will become applicable. Providers and deployers of AI systems need to learn how to apply the different articles of the act, while notified bodies and certification authorities will need to establish a process to assess these systems according to the regulations. Many standardization efforts are already underway, but there will be significant uncertainty about how the AI Act has to be applied and how it relates to different aspects of AI currently under research.
Regulations such as the AI Act are especially important in high-risk artificial intelligence (AI) systems to ensure that they are safe and respect fundamental rights and values (van_kolfschooten_eu_2024). The AI Act follows a risk-based approach: the higher the risk, the stricter the rule. Especially in high-risk systems, it describes many obligations for the provider and the deployer of AI systems.
@@
cdss, which may fall under the definition of a high-risk AI system in the EU AI Act, place the responsibility for the high-risk decision on the end-user (typically, clinicians). However, the AI Act does not directly address end-user needs and instead primarily focuses on regulating the providers and deployers of AI systems. One of the few mentions of end-users is the statement that high-risk AI systems must be designed so that humans can effectively oversee these systems. However, there are no technical requirements that prioritize the user’s needs at the center of its goals (van_kolfschooten_eu_2024).
In contrast, nowadays more explainable artificial intelligence (XAI) researchers suggest putting end-users and their needs at the center of attention (langer_what_2021). XAI methods provide descriptions of the outputs of opaque AI models, thereby satisfying the needs of the end-users. This is very helpful in high-risk applications such as healthcare. Previous work has started exploring the role that XAI plays w.r.t. ensuring the trustworthiness of high-risk applications under the EU AI Act (panigutti_role_2023; kiseleva_transparency_2022; longo_explainable_2024; esmaeilzadeh_challenges_2024; bibalLegalRequirementsExplainability2021). XAI can contribute to implementing transparency and human oversight, which are key principles of the AI Act (panigutti_role_2023). Taking this line of thinking further, we believe that XAI—if designed correctly—could play a crucial role in fulfilling the obligations of deployers and providers of AI systems. However, we acknowledge that XAI alone cannot fulfill all obligations under the AI Act. Full compliance requires additional measures such as the development of harmonized standards, risk management frameworks, and legal assessments, which are beyond the scope of XAI methods. Therefore, we ask:
@@
Research on the AI Act and research on XAI are often conducted independently, lacking interdisciplinary collaboration (bexAILawTransdisciplinary2025; gal_bridging_2023). By providing a joint understanding of the role of XAI in the AI Act from an interdisciplinary perspective representing various stakeholder needs, we…
@@
    provide an analysis of the implications of the AI Act in relation to adopting AI in healthcare (see Section˜4)
@@
    describe how the needs of the CDSS stakeholders can be captured through XAI (see Section˜5)
@@
    assess the similarities and contrasts between the legal requirements of the AI Act and the stakeholder desiderata of XAI (see Section˜6).
With this, our work is the first that analyzes the role of XAI in the AI Act in a cross-functional team, on a real-world CDSS that takes user-centered needs into account and contributes to the state-of-the-art by mapping obligations of the EU AI Act with the obligations of stakeholders and how to fulfill them using XAI. With this, we encourage researchers and practitioners from the XAI community to reflect on their role in relation to the AI Act by bridging the gap with end-users.
@@ 2 Related Work and Background
In this section, we give an overview of the AI Act and XAI. We also discuss relevant related work. Several studies have already addressed the question of how AI can be applied in healthcare, incorporating ethical and legal considerations (aboy_navigating_2024; porsdam_mann_eu_2024; karimian_ethical_2022; kiseleva_transparency_2022). The function of XAI in fostering transparency and human oversight has also been the subject of considerable debate (panigutti_role_2023). Other works have dealt with the role of XAI in the AI Act (e.g., (gal_bridging_2023)) and have given an overview of the AI Act and XAI (e.g., (panigutti_role_2023)). Nevertheless, it remains unexplored how XAI can be effectively utilized to fulfill the obligations of stakeholders as defined in the AI Act.
@@ 2.1 Background: The EU AI Act
In 2024, the European Union ratified the regulation of AI as technology (aiact). The new regulation was motivated by a need to create trust for AI with respect to product safety, health, and fundamental rights. The AI Act thus incorporates AI systems into the EU’s existing framework for product safety. Certain AI systems will need to undergo conformity assessment and subsequently carry the CE-mark as a sign of being a certified product, just like elevators, machinery, and medical devices (see Section˜2.1).
The regulation also gives authorities an important role in guaranteeing the responsible development and usage of AI, as public sector bodies are to conduct market surveillance but also foster AI innovation (see e.g., (regulatingtrust)). There are two central actors from the perspective of the regulation: the provider (see Section˜2.1) and the deployer (see Section˜2.1) of an AI system. These actors are obliged to take on responsibility for product safety.
@@
The conformity assessment for CE-marking AI systems can then be done in two ways. The first is through self-assessment, where the provider of the system certifies that the requirements are met. This is also referred to as internal conformity assessment. The second conformity assessment procedure involves a notified body, an organization that has been accredited to be competent in conformity assessment of certain products. The role of the notified body is to give a neutral assessment of the extent to which the requirements are met. The nature of the AI system will determine which process is to be applied (conformityassessment).
@@
There will be harmonized standards issued by the European standardization organizations CEN and CENELEC that further detail the requirements in Art. 9 to 15 and 17. An overview of the standardization effort can be found on the two organizations’ joint technical committee’s website (europeancommitteeforstandardizationCENCENCLC2025).
@@ 2.2 Background: XAI Terminology
As indicated, current research on XAI provides a new perspective on the relationship between an AI system and its affected parties. Therefore, we provide a brief nomenclature of relevant key concepts and terms. We define the key terms related to XAI, examine the various stakeholder terminologies, and clarify a common terminology of XAI [^1] We acknowledge that terms like interpretability, explainability, transparency, or trust are used differently across disciplines. In this paper, we adopt definitions from the technical XAI literature to guide system design and evaluation. Our focus is on their operational meaning in the context of AI Act compliance, rather than engaging in broader philosophical debates. methods.
@@
The terms explainability, interpretability, and transparency are often used interchangeably (al-ansari_user-centered_2024; renftle_what_2024; hoffman_measures_2023; barredo_arrieta_explainable_2020; mohseni_multidisciplinary_2021). We distinguish these terms by categorizing them into approaches like Explainable AI and Interpretable AI, and desired properties such as Interpretability, Transparency, and Explainability (mohseni_multidisciplinary_2021). Explainable AI provides human-understandable explanations for predictions made by complex AI systems that are not inherently interpretable (wang_designing_2019). Interpretable AI, on the other hand, requires no additional methods for human understanding, as these systems are inherently interpretable (mohseni_multidisciplinary_2021). Interpretability refers to the ability to explain or provide meaning in understandable terms, while Transparency indicates a model is interpretable by design (barredo_arrieta_explainable_2020).
@@
XAI goals such as trustworthiness, informativeness, and fairness have been widely studied (al-ansari_user-centered_2024; renftle_what_2024; hoffman_measures_2023; barredo_arrieta_explainable_2020; mohseni_multidisciplinary_2021). Recently, these goals have shifted to focus more on specific stakeholders (samek_towards_2019; biecek2024position), their desiderata (langer_what_2021), and expertise levels (mohseni_multidisciplinary_2021). Stakeholders’ interests, goals, expectations, needs, and demands regarding AI systems are referred to as their desiderata (langer_what_2021).
Understanding different stakeholders in XAI is essential for creating systems that meet various desiderata (langer_what_2021). Different groups have distinct interests, and recognizing these differences is key to developing methods that align with their specific desiderata. The success of an XAI method relies on satisfying stakeholders’ desiderata, which drives explainability approaches that enhance understanding. langer_what_2021, following barredo_arrieta_explainable_2020, categorize stakeholders into five classes: users, (system) developers, affected parties, deployers, and regulators. This classification ensures all relevant needs are considered, with the challenge being to provide suitable explanations for each group (mohseni_multidisciplinary_2021).
Following mohseni_multidisciplinary_2021, end-users are divided into three groups based on their expertise level: Novice Users, who use AI systems daily with little machine learning (ML) knowledge; Data Experts, including data scientists and domain experts using ML for analysis; and AI Experts, comprising ML scientists and engineers designing and implementing ML algorithms and XAI methods.
@@
We adopt the taxonomy proposed by speith_review_2022 as it provides a structured and widely used framework for categorizing XAI methods. These categories are later applied in our methodology (see Section˜5.2) to guide the selection of suitable XAI techniques for our CDSS.
@@
Scope refers to local explanations that detail specific input-output relations, while global explanations describe the overall ML model. Examples include model visualizations (liu_towards_2017) and decision rules (lakkaraju_interpretable_2016).
Other taxonomies include input data (e.g., image, time series), output format (e.g., numerical, visual), and problem type (e.g., classification, regression).
@@ 2.3 Related Work
As previously mentioned, several studies address the question of how AI can be applied in healthcare in general. Table˜1 provides an overview of the related work papers and indicates how we position our own contribution.
@@
Paper
AI Act
XAI
Healthcare
Real World
User-Centric
gal_bridging_2023
pavlidis_unlocking_2024
fresz_how_2024
onitiu_limits_2023
solaiman_regulating_2024
van_kolfschooten_eu_2024
panigutti_role_2023
@@
gal_bridging_2023 explore the transparency gap between the AI Act and XAI understanding, emphasizing the differing conceptualizations. They focus on stakeholder engagement and how XAI can address these concerns, aligning with our approach. Identified solutions include concentrating on legal definitions, AI system output, and data-related requirements to meet transparency standards. It is important to note that transparency requirements are broad and should not be seen as concrete engineering instructions. XAI contributes to transparency but should not be viewed as a standalone solution; instead, it facilitates transparency while adhering to its foundational principles.
@@
pavlidis_unlocking_2024 explores the role of XAI as a tool to bridge the gap between the opaque nature of ML decisions and the human demand for comprehensibility. In essence, XAI serves as a mechanism to align the objectives of the AI Act, such as enhancing transparency, with the preferences of end-users. The author examines the integration of XAI into the AI Act, with a focus on critical issues like standard setting, oversight, and enforcement. fresz_how_2024 review legal requirements for explainability of AI systems from the AI Act. Furthermore, they describe legal bases for the use of XAI and derive requirements that the XAI methods need to fulfill. All three articles deal with the role of XAI in the AI Act. Our work extends this view by elaborating the role of XAI in a real-world healthcare scenario and by incorporating the end-user needs.
@@
A more domain-specific approach is presented by onitiu_limits_2023, who aims to translate technical safeguards in Art. 13-14 of the AI Act proposal for the use of medical diagnostic systems. The author proposes a broader perspective on the socio-legal implications of technical safeguards, such as post-hoc explainability for decision-making. onitiu_limits_2023 emphasizes that post-hoc XAI methods must focus on harmonizing human expertise with patient-oriented values.
@@
Similarly, solaiman_regulating_2024 explore the AI Act in the healthcare domain with a user-centric focus on AI. Hereby, XAI plays a subordinate role. Instead, the authors focus on the doctor-patient relationship that evolved from a relation where the doctors held the dominant role and the patient follows their guidance without explanation, to a relationship where the patient is seen as an active participant in their healthcare decisions. AI and especially XAI can have an important impact providing more information to both clinicians and patients, thus enhancing patient autonomy.
@@
van_kolfschooten_eu_2024 analyze the implications of the AI Act for the healthcare domain. Here, they give an overview of the obligations for various healthcare stakeholders. Then they use their insights to provide a guideline to address the unique needs of the stakeholders and make recommendations for the implementation and standardization of the act. They also highlight the importance of collaborating with different stakeholders like AI developers, healthcare professionals, patient communities, and regulators to develop those guidelines.
@@
In a more application-oriented study, panigutti_role_2023 conduct an interdisciplinary analysis of the AI Act’s approach to opaque AI systems, aiming to clarify XAI’s role. They focus on Art. 13, concerning transparency, and Art. 14, regarding human oversight, providing an overview of XAI terminology and key concepts. The authors also review the XAI field, noting challenges and limitations consistent with previous findings. They analyze how the AI Act addresses AI system opacity through transparency and human oversight, presenting a use case involving an AI-based proctoring system without XAI tools. In contrast, our study extends beyond Art. 13 and 14 and examines other articles that establish obligations for healthcare stakeholders.